A simple random sample of 25,000 individuals are surveyed in order to determine the prevalence of individuals that contracted the flu in the past year. Assuming 4,250 individuals indicated they had been diagnosed with the flu at some point in the past year, what is the prevalence of flu for the past year, as indicated from the survey participants?

A. 8.30%

B. 83.0%

C. 17.0%

D. 1.70%
Answer #1

The correct option is C. 17.0%.

Total number of random sample N = 25,000

Individuals affected with flu n = 4,250

Then prevalence of flu = n / N = 4250/25000 = 0.17 = 17 %.